OpenStudy (anonymous):

if f(x)=|x-1| and g (x) = 5x -4, then which of the following is false? (A) fog (x) =|5x-5| (B) gof (x) = 5|x-1| -4 (C) gog (1) = 4 (D) fof (-1) = 1

OpenStudy (anonymous):

i got -4 for the c) part

OpenStudy (anonymous):

Hi Julian....do u think (C) is right ans to the question, i'm confused

OpenStudy (anonymous):

well make the operation gog (1) and see what u got do u have problems finding gog (1)?

OpenStudy (anonymous):

it's been a long long time i did maths? suddenly for computer programming it is required.....Thanx for ur help frnd....and did u check other options?

OpenStudy (anonymous):

If f(x)= 5-x,then which of the following is false ? (A) is strictly increasing in (0, (B) is strictly decreasing in (0, (C) is strictly decreasing in (- 0) (D) =R(real numbers) =(0,

OpenStudy (anonymous):

well maths is requeried even for programming u have there some math operators i think that the logic in math and programming is very similar i check the oither options

OpenStudy (anonymous):

increasing in (0, ?????? that is not hte complete interval

OpenStudy (anonymous):

that is alpha. In A) (0,alpha) B) (0,alpha) C) ( -alpha,0) D) (0,alpha)

OpenStudy (anonymous):

alpha? well i cant give an answer without know the value of "alpha" that function is decreasing in its entire domain

OpenStudy (anonymous):

alphs=Infinite

OpenStudy (anonymous):

that function is decreasing on its entire domain

OpenStudy (anonymous):

that Means (B) Or (C) ?

OpenStudy (anonymous):

R u there Julian?

OpenStudy (anonymous):

If f(x)=|x+2| then, (A) f(x+y)=f(x)+f(y) (B) f|x|=|f(x)| (C) f(xy)=f(x).f(y) (D) none of these